Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2024 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 93     Entry time: 02 May 2024, 23:19
Author: Juncheng E 
Group: SPB 
Subject: Shift summary 

X-ray delivery

  • 9.3 keV, 3.3 mJ, 190 pulses/train

Optical later delivery

  • For jet illumination

Achievements/Observations

  • Aligned NKB mirrors, slits again
  • Removed the AGIPD shield
  • Moved slit3 and slit4 to a fresh position after damage

Issues

  • Litframe finder blocked collecting dark images for AGIPD. The problem was solved after resetting the Litframe finder.
  • Slit3 and slit4 were damaged
ELOG V3.1.4-7c3fd00